# Are Helium Atmospheres Hiding as Volatile-rich in Current Exoplanet Analysis Frameworks?

### v1

@@ 4.3.2 Outflow Fractionation
Another promising avenue to help disentangle the H2- and He-dominated scenarios is through the fractionation of their outflow. Outflow observations mostly trace the upper atmosphere, yet the fractionation is controlled by the coupling in the lower atmosphere (Xing et al., 2023).
@@
Numerous giant exoplanets, incl. HD 209458 b, present an outflow with a subsolar He/H ratio (e.g., Lampón et al., 2020, 2021; Khodachenko et al., 2021; Shaikhislamov et al., 2021). While these subsolar He/H outflow ratio were originally linked to a low atmospheric abundance of helium, Xing et al. (2023) recently showed using a self-consistent multi-fluid hydrodynamic model that He-mass fractionation can explain the outflow of HD 209458 b. In other words, HD 209458 b does not need to be He-poor to explain the low He/H ratio of its outflow. In fact, the opposite can be true as it can preferentially lose hydrogen over billions of years.
4.3.3 Atmospheric Chemistry
@@
A third avenue to considering leveraging to help disentangle the H2- and He-dominated scenarios relates to the different chemical composition expected for each of them. One of the most notable differences/markers expected relates to CO and CO2 being the main reservoir of carbon as CH4 will have to compete with H2O for the (relatively) scarce hydrogen (Hu et al., 2015; Guzmán-Mesa et al., 2022). The four molecules above being strong absorbers, their relative abundances may provide the most constraining means to disentangle between the two atmospheric scenarios–with the caveat that the interpretation of the abundances will be dependent (i.e., sensitive) to the atmospheric/chemistry models.
5 Conclusion
@@
By freeing the helium-to-hydrogen ratio in atmospheric retrievals of HD 209458 b, we show that the surprisingly-large water abundance reported in the literature can disappear once the scale height is allowed to decrease through helium enrichment—rather than through an artificially high water content (or, more broadly, metallicity). This provides a new avenue for interpreting JWST spectra and upcoming composition/metallicity trends. It is also a reminder that as data quality increases, hidden assumptions in analysis frameworks may exacerbate biases in our inferences.
@@ Acknowledgements
The authors thank C. Cherubim, I. Gordon, R. Hargreaves, S. Seager, L. Wiesenfeld, R. Wordworth, and Q. Xue for valuable inputs. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the present work. J.d.W. and P.N. acknowledge support from the European Research Council (ERC) Synergy Grant under the European Union’s Horizon 2020 research and innovation program (grant No. 101118581 — project REVEAL).

### v2

@@ IV.3.2 Outflow Fractionation
Another promising avenue to help disentangle the H2- and He-dominated scenarios is through the fractionation of their outflow. Outflow observations mostly trace the upper atmosphere, yet the fractionation is controlled by the coupling in the lower atmosphere (Xing2023).
@@
Numerous giant exoplanets, incl. HD 209458 b, present an outflow with a subsolar He/H ratio (e.g., Lampon2020a; Lampom2021; Khodachenko2021; Shaikhislamov2021). While these subsolar He/H outflow ratio were originally linked to a low atmospheric abundance of helium, Xing2023 recently showed using a self-consistent multi-fluid hydrodynamic model that He-mass fractionation can explain the outflow of HD 209458 b. In other words, HD 209458 b does not need to be He-poor to explain the low He/H ratio of its outflow. In fact, the opposite can be true as it can preferentially lose hydrogen over billions of years.
IV.3.3 Atmospheric Chemistry
@@
A third avenue to considering leveraging to help disentangle the H2- and He-dominated scenarios relates to the different chemical composition expected for each of them. One of the most notable differences/markers expected relates to CO and CO2 being the main reservoir of carbon as CH4 will have to compete with H2O for the (relatively) scarce hydrogen (Hu2015; Guzman2022). The four molecules above being strong absorbers, their relative abundances may provide the most constraining means to disentangle between the two atmospheric scenarios–with the caveat that the interpretation of the abundances will be dependent (i.e., sensitive) to the atmospheric/chemistry models.
IV.4 Other Considerations for Robust Inferences
While no exoplanet spectra has been interpreted as sign of an unexpectedly-high metallicity (e.g., water abundance) to date, a substantial fraction of interpretations invoke clouds (see, e.g., Espinoza2025, and references therein). In addition, the abundances reported in some instances appear biased to low values due to priors. For example, while Bell2023 reports a () for WASP-80 b their PPD hits the upper bound of their priors well within 1 (see their Extended Data Fig. 3) meaning that without such hard edge the favored water abundance would likely be closer to . Jointly, these facts call for (re)analyses with a special care for the complex dependencies discussed here.
V Conclusion
@@
By freeing the helium-to-hydrogen ratio in atmospheric retrievals, we show that helium enrichment and high water content (or, more broadly, metallicity) can explain the data at hand—in addition to clouds. This provides a new avenue for interpreting JWST spectra and upcoming composition/metallicity trends. It is also a reminder that as data quality increases, hidden assumptions in analysis frameworks may exacerbate biases in our inferences.
@@ Acknowledgements
The authors thank the Reviewer, C. Cherubim, J. Fortney, I. Gordon, R. Hargreaves, Z. Lin, S. Seager, D. Thorngren, L. Wiesenfeld, R. Wordworth, and Q. Xue for valuable inputs. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the present work. J.d.W. and P.N. acknowledge support from the European Research Council (ERC) Synergy Grant under the European Union’s Horizon 2020 research and innovation program (grant No. 101118581 — project REVEAL).